# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v1

@@ Abstract
Fine-tuning large language models (LLMs) with classic first-order optimizers entails prohibitive GPU memory due to the backpropagation process. Recent works have turned to zeroth-order optimizers for fine-tuning, which save substantial memory by using two forward passes. However, these optimizers are plagued by the heterogeneity of parameter curvatures across different dimensions. In this work, we propose HiZOO, a diagonal Hessian informed zeroth-order optimizer which is the first work to leverage the diagonal Hessian to enhance zeroth-order optimizer for fine-tuning LLMs. What’s more, HiZOO avoids the expensive memory cost and only increases one forward pass per step. Extensive experiments on various models (350Msimilar-$\sim$66B parameters) indicate that HiZOO improves model convergence, significantly reducing training steps and effectively enhancing model accuracy. Moreover, we visualize the optimization trajectories of HiZOO on test functions, illustrating its effectiveness in handling heterogeneous curvatures. Lastly, we provide theoretical proofs of convergence for HiZOO. Code is publicly available at https://anonymous.4open.science/r/HiZOO-27F8.
@@ 1 Introduction
Figure 1: The toy example with heterogeneous curvatures, where $X$ is the sharp dimension and $Y$ is the flat dimension. HiZOO updates along the direction with greater curvature and converges obviously more quickly than MeZO. The corresponding loss curves are in Section 5.6.
Fine-tuning pre-trained LLMs is necessary for specific downstream tasks and first-order optimizers are widely used, such as Adam (Kingma & Ba, 2015) and AdamW (Loshchilov & Hutter, 2019). Nevertheless, these optimizers are accompanied by substantial memory consumption, primarily due to the inherent backpropagation process. To address these limitations, MeZO (Malladi et al., 2023) proposed to utilize a memory-efficient zeroth-order optimizer to estimate the gradient with just two forward passes per step, achieving numerous memory reduction.
However, loss functions of LLMs often exhibit heterogeneous curvatures across different parameter dimensions, as documented in recent studies (Sagun et al., 2017; Ghorbani et al., 2019; Zhang et al., 2020). This significant difference in curvature can lead to instability or decelerated training, as shown in Figure 1. The Hessian can be leveraged in optimizer to effectively adjust the magnitude of the parameter updates (Liu & Li, 2023; Yao et al., 2021; Anil et al., 2021), solving the above dilemma. Unfortunately, in the context of zeroth-order optimization, one cannot directly compute the Hessian atop first-order derivatives.
In light of above, we propose HiZOO, leveraging the diagonal Hessian in zeroth-order optimizer to improve the convergence when training the model with heterogeneous curvatures. As shown in Figure 2, HiZOO can significantly reduce the training steps and improve the accuracy. Here we summarize our key contributions as follows:
Figure 2: HiZOO achieves up to 30×30$\times$30 × speedup over MeZO while achieving better accuracy. (a) With few-shot setting ($K=16$), HiZOO achieves 93.2% while MeZO achieves 90.5%. (b) HiZOO achieves 69.6% while MeZO achieves 66.1%. (c) HiZOO achieves 60.6% while MeZO achieves 57.7%. All the comparative experiments are conducted with the same parameters shown in Appendix D.1.
@@
    In HiZOO, we estimate the Hessian in zeroth-order optimization to fine-tune LLMS for the first time. Compared with classic zeroth-order optimizers, HiZOO merely increases one more forward pass per step. What’s more, HiZOO utilizes diagonal Hessian and reduces the corresponding memory cost from $\mathcal{O}(d^{2})$ to $\mathcal{O}(d)$.
@@
    We provide theoretical analysis to prove that HiZOO provides an unbiased estimation of the Hessian and successfully achieve global convergence in even cases with heterogeneous curvatures across parameter dimensions.
@@
    We conduct extensive experiments across model types (masked language models and auto-regressive language models), model scales (from 350M to 66B), and downstream tasks (classification, multiple-choice, and generation) to verify the effect of the HiZOO. Especially on SST-2 dataset, HiZOO can reduce the training time by up to 30 times (from 100k steps to 3k steps), achieving the same performance. What’s more, HiZOO finally improves the accuracy by 3% on SST-2, even better than standard full-parameter fine-tuning. When tuning OPT-66B, HiZOO reduces the model training time and performs better with up to 5% improvement and 2% improvement on average.
@@
    We demonstrate HiZOO’s compatibility with full-parameter tuning, LoRA and prefix-tuning in Section 5. Specifically, HiZOO (prefix) outperforms MeZO (prefix) on most of the datasets and tasks across all model scales.
  5. 5.
    Further exploration in Section 5.3 showcases that HiZOO can achieve better performance in optimizing non-differentiable objectives such as F1 score. Using the same training steps, HiZOO significantly outperforms the MeZO ’s results with 6.5% on average.
@@ 2 Related Works
Here we present a concise summary of related works on optimizers used in fine-tuning LLMs. For a more detailed review, please refer to Appendix A.
2.1 First-Order Optimizer in LLMs
First-order optimizers, such as Gradient Descent (GD), Momentum, Adagrad (Duchi et al., 2011), are fundamental in many areas like computer vision, natural languagle processing (NLP). Among them, Adam (Kingma & Ba, 2015) plays a dominant role due to its fast convergence and is often chosen for training and fine-tuning LLMs. AdamW (Loshchilov & Hutter, 2019) improves upon Adam by adding the weight decay to alleviate overfitting.
2.2 First-Order Optimizer Enhanced with Hessian
On the other hand, researchers incorporated second-order information (Hessian) to provide richer guidance for gradient descent during the training. For example BROYDEN (1970) , Nesterov & Polyak (2006) and Conn et al. (2000) utilized curvature information to pre-condition the gradient; Magoulas et al. (1999) applied diagonal Hessian as the pre-conditioner; Martens (2010) approximated the Hessian with conjugate gradient. Sophia (Liu & Li, 2023) used a light-weight estimate of the diagonal Hessian for pre-training LLMs. Despite their potential, above optimizers require the enormous GPU-memory cost.
2.3 Zeroth-Order Optimizer
Zeroth-order optimizers, with just forward passes, can greatly reduce the memory consumption. Methods like SPSA (Spall, 1992) have been shown to perform well in non-convex multi-agent optimization (Tang et al., 2021; Hajinezhad & Zavlanos, 2018) or generating black-box adversarial examples (Chen et al., 2017; Cai et al., 2021; Liu et al., 2019a; Ye et al., 2019). Recently, MeZO (Malladi et al., 2023) firstly adapted the classical ZO-SGD method to fine-tune LLMs, achieving comparable performance with significant memory reduction.
@@ 3 Methods
In the following, we briefly introduce the classical ZO gradient estimator SPSA (Spall, 1992), which is used in MeZO. Then we describe how HiZOO estimates diagonal Hessian and cooperates with zeroth-order optimizer.
@@ 3.1 Preliminaries
$g_{\mu}(\theta)=\frac{\mathcal{L}(\theta+\mu u;\mathcal{B})-\mathcal{L}(\theta-\mu u;\mathcal{B})}{2\mu}u\approx$ $uu^{\top}\nabla\mathcal{L}(\theta;\mathcal{B}),$
@@ 3.2 Diagonal Hessian Estimator
Storing the exact full spectral Hessian ($d\times d$) requires $\mathcal{O}(d^{2})$ memory (Yao et al., 2018; Xu et al., 2019; Dembo et al., 1982), which is sufficient but never necessary. In HiZOO, we just estimate and retain only the diagonal Hessian which requires $\mathcal{O}(d)$ memory. It serves as a pre-conditioner to scale the direction and magnitude of the model parameter updates according to their respective curvatures.
Drawing from the lemma presented in MiNES (Ye, 2023):
@@
where $H$ is the Hessian $\nabla^{2}\mathcal{L}(\theta)$ and $\Sigma$ is the approximate Hessian inverse matrix with $\Sigma^{-1}\approx\nabla^{2}\mathcal{L}(\theta)$.
Thus, we can approximate the diagonal Hessian by the zeroth order oracles. Firstly, we will access to the $\mathcal{L}(\theta+\mu\Sigma^{1/2}u)$, $\mathcal{L}(\theta-\mu\Sigma^{1/2}u)$ and $\mathcal{L}(\theta)$. Through the Taylor’s expansion, we yield the following results:
$\mathcal{L}(\theta+\mu\Sigma^{1/2}u)=\mathcal{L}(\theta)+\mu\langle\mathcal{L}(\theta),\Sigma^{1/2}u\rangle$ (2)
$+\frac{\mu^{2}}{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta)\Sigma^{1/2}u+\alpha(\theta,\mu\Sigma^{1/2}u).$
@@
$\mathcal{L}(\theta-\mu\Sigma^{1/2}u)=\mathcal{L}(\theta)-\mu\langle\mathcal{L}(\theta),\Sigma^{1/2}u\rangle$
$+\frac{\mu^{2}}{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta)\Sigma^{1/2}u+\alpha(\theta,-\mu\Sigma^{1/2}u).$
Then we can calculate the $\Delta\mathcal{L}$ by:
$\Delta\mathcal{L}=$ $\mathcal{L}(\theta+\mu\Sigma^{1/2}u)+\mathcal{L}(\theta-\mu\Sigma^{1/2}u)-2\mathcal{L}(\theta)$
$=$ $\mu^{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta)\Sigma^{1/2}u+\alpha(\theta,\mu\Sigma^{1/2}u)$
$+\alpha(\theta,-\mu\Sigma^{1/2}u).$
@@
$\frac{\Delta\mathcal{L}}{\mu^{2}}=u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta)\Sigma^{1/2}u+\mathcal{O}(\mu).$
@@
$\frac{1}{2}\mathbb{E}\bigg{[}\frac{\Delta\mathcal{L}}{\mu^{2}}\cdot\left(\Sigma^{-1/2}uu^{\top}\Sigma^{-1/2}-\Sigma^{-1}\right)\bigg{]}=\nabla^{2}\mathcal{L}(\theta)+\mathcal{O}(\mu).$
@@
Above equation shows that we can approximate the diagonal entries of $\nabla^{2}\mathcal{L}(\theta)$ by $\mathrm{diag}(\Sigma_{t}^{\prime}(\theta))$, requiring just one more forward pass per step compared with MeZO.
Due to the presence of noise in the calculation of the Hessian, we utilize exponential moving average (EMA) to denoise the diagonal Hessian estimation, which commonly used in moments of gradients of Adam.
$\Sigma_{t+1}=(1-\alpha_{t})\Sigma_{t}+\alpha_{t}\left|\mathrm{diag}(\Sigma^{\prime}_{t})\right|.$ (4)

### v3

@@ Abstract
Fine-tuning large language models (LLMs) is necessary for specific downstream tasks, but the classic adaptive first-order optimizer entails prohibitive GPU memory because of backpropagation. Recent works such as MeZO have turned to zeroth-order optimizers for fine-tuning, which reduce substantial memory by using just two forward passes. However, heterogeneous curvatures across different parameter dimensions in LLMs often cause convergence instability or even failure. In this work, we propose HiZOO, a diagonal Hessian informed Zeroth-Order Optimizer , which is the first to leverage the diagonal Hessian to enhance ZOO for fine-tuning LLMs. We provide the theoretical proof for HiZOO and visualize the optimization trajectories on the test functions. Extensive experiments on various models (RoBERTa, OPT, Phi-2, and LLama3, with 350Msimilar-$\sim$66B parameters) indicate that HiZOO significantly reduces the number of training steps and improves model accuracy. For example, on the SST2 task, HiZOO achieves an 8x speed-up and better accuracy. Even when scaled to 66B-model, HiZOO outperforms MeZO with up to 5.1% absolute improvement. We also propose HiZOO-L, which reduces the Hessian memory cost to 10% of the MeZO, while maintaining almost same performance. Compared with ZO-Adam, HiZOO-L achieves a 4.3% absolute improvement, just using 50% of the GPU memory. Code is available at https://anonymous.4open.science/r/HiZOO-27F8.
Equal contribution. Correspondence to: yehaishan@xjtu.edu.cn
@@ 1 Introduction
Figure 1: (Left) Comparison of HiZOO, MeZO and Adam. (Right) The test example with heterogeneous curvatures. HiZOO updates along the direction with greater curvature ($X$) and converges more quickly than MeZO. The corresponding loss curves are shown in Section 3.5.
Fine-tuning pre-trained large language models (LLMs) for specific tasks has garnered significant attention. As the number of model parameter increases, full parameter fine-tuning (FT) becomes markedly memory-intensive. To alleviate the memory limitations, Parameter Efficient Fine-Tuning (PEFT) methods [19, 24, 10, 60, 35] have been developed, which only fine-tune a small number of (extra) model parameters. As a result, they significantly reduce computational and storage costs, while achieving performance comparable to a fully fine-tuned model.
First-order optimizers such as Adam [22] and AdamW [29] are widely used to fine-tune LLMs. Nevertheless, using these optimizers still leads to substantial memory consumption, primarily due to the inherent backpropagation process. To address these limitations, MeZO [32] proposed to utilize a zeroth-order optimizer (ZOO) to estimate the gradient with just two forward passes per step, no need for backpropagation anymore. This achieves numerous memory reduction and makes it accessible to train and store LLMs on consumer hardware.
However, parameters of LLMs often exhibit heterogeneous curvatures across different dimensions, as documented in recent studies [39, 16, 55]. This significant difference in curvature would make model converge towards the saddle point, slowing down the convergence speed, as shown in Figure 1 (right). Since the incorporation of Hessian to measure the curvature properties of the loss landscape, second order methods [26, 50, 2] can solve this sub-optimal behaviour. Unfortunately, in the context of zeroth-order optimization, one cannot directly compute the Hessian atop first-order derivatives.
In light of above, we propose HiZOO, as shown in Figure 1 (left), which estimates the diagonal Hessian by one more forward pass. HiZOO can act as a pre-conditioner, directly adjusting the update size of different parameters according to their curvatures. So that it can improve the model convergence when encountered with heterogeneous curvatures. As shown in Figure 2, HiZOO can significantly reduce the training steps and improve the accuracy. Here we summarize our key contributions as follows:
Figure 2: Performance of MeZO, HiZOO and HiZOO-L on SST2 task, when fine-tuning RoBERTa-large, OPT-13B, Llama3(8B) models. HiZOO can achieve 8$\times$ speedup and 1.55% absolute accuracy improvement compared with MeZO. Experiments are conducted with the same learning rate.
@@
    In this work, we estimate the Hessian in zeroth-order optimizer to fine-tune LLMs for the first time. Our HiZOO reduces the total number of forward passes required for model convergence and achieves better accuracy. By utilizing diagonal Hessian, HiZOO reduces the corresponding memory cost from $\mathcal{O}(d^{2})$ to $\mathcal{O}(d)$. Furthermore, we propose HiZOO-L, reducing the memory usage of Hessian to 10% of the MeZO.
@@
    We provide theoretical analysis to prove that HiZOO provides an unbiased estimation of the Hessian. Also, we illustrate how HiZOO utilizes Hessian to improve the convergence process by visualizing the optimization trajectories on test functions.
@@
    We conduct extensive experiments across different models (RoBERTa-large, OPT, Llama3 and Phi-2) with scales from 350M to 66B, different methods (FT, LoRA, prefix), and different downstream tasks (classification, multiple-choice, and generation) to verify the effect of the HiZOO. For example, on SST2 task HiZOO achieves a 1.55% accuracy improvement and 8×8$\times$8 × speedup over MeZO on average across different models. Even on OPT-66B, HiZOO outperforms better than MeZO with up to 5% improvement.
@@
    Further exploration in Section 4.3 showcases that HiZOO can achieve better performance in optimizing non-differentiable objectives such as F1 score. Specifically, HiZOO significantly outperforms the MeZO ’s results with 6.5% on average.
@@ 2 Related Works
Here we present a concise overview on optimizers used in fine-tuning LLMs(details in Appendix A).
First-Order Optimizer in LLMs    First-order optimizers, such as Gradient Descent (GD), Momentum, Adagrad [12], are fundamental in many areas like computer vision, natural languagle processing (NLP). Among them, Adam [22] plays a dominant role due to its fast convergence and is often chosen for training and fine-tuning LLMs. AdamW [29] improves upon Adam by adding the weight decay to alleviate overfitting. But both of them requires lots of memory cost due to the backpropagation process.
First-Order Optimizer Enhanced with Hessian    On the other hand, researchers incorporated second-order information (Hessian) to provide richer guidance for gradient descent during the training. For example BROYDEN [4] , Nesterov & Polyak [34] and Conn et al. [7] utilized curvature information to pre-condition the gradient; Magoulas et al. [31] applied diagonal Hessian as the pre-conditioner; Martens [33] approximated the Hessian with conjugate gradient. Sophia [26] used a light-weight estimate of the diagonal Hessian for pre-training LLMs. Despite their potential, above optimizers require the enormous GPU-memory cost.
Zeroth-Order Optimizer    Zeroth-order optimizers, with just forward passes, can greatly reduce the memory consumption. Methods like SPSA [42] have been shown to perform well in non-convex multi-agent optimization [44, 18] or generating black-box adversarial examples [6, 5, 27, 52]. Recently, MeZO [32] firstly adapted the classical ZO-SGD method to fine-tune LLMs, achieving comparable performance with significant memory reduction. But these methods often struggle with heterogeneous curvatures.
@@ 3 Methods
In the following, we briefly introduce the classical ZO gradient estimator SPSA [42], which is used in MeZO. Then we describe how HiZOO estimates diagonal Hessian and cooperates with ZOO .
@@ 3.1 Preliminaries
$g^{\prime}_{\mu}(\theta_{t})=\frac{\mathcal{L}(\theta_{t}+\mu u;\mathcal{B})-\mathcal{L}(\theta_{t}-\mu u;\mathcal{B})}{2\mu}u\approx$ $uu^{\top}\nabla\mathcal{L}(\theta_{t};\mathcal{B}),$
@@
3.2 Hessian Informed Zeroth-Order Optimization
1:parameters $\theta\in\mathbb{R}^{d}$, loss $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, learning rate schedule ${\eta_{t}}$, smooth scale $\alpha_{t}$, diagonal Hessian $\Sigma_{0}$
2:for $t=1,...,T$ do
3:     Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
4:     $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
5:     $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
6:     $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
7:     $\theta$leftarrow PerturbParameters($\theta$, $-2\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
8:     $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
9:     $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$) $\triangleright$ Reset parameters before descent
10:     $\Sigma^{\prime}_{t}=\frac{1}{2\mu^{2}}(\ell_{+}+\ell_{-}-2\ell)(\Sigma^{-1/2}_{t-1}u_{i}u_{i}^{\top}\Sigma^{-1/2}_{t-1})$triangleright Update diagonal Hessian
11:     $\Sigma_{t}^{-1}=(1-\alpha_{t})\Sigma_{t-1}^{-1}+\alpha_{t}\left|diag(\Sigma^{\prime}_{t})\right|$
12:     projected_grad $\leftarrow(\ell_{+}-\ell_{-})\ast\Sigma^{1/2}_{t}/2\mu$
13:     Reset random number generator with seed $s$triangleright For sampling $u_{i}$
14:     for $\theta_{i}\in\theta$ do
15:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
16:         $\theta_{i}\leftarrow\theta_{i}-\eta_{t}\ast$ projected_grad $\ast u_{i}$
17:     end for
18:end for
19:
20:function PerturbParameter($\theta$, $\mu$, $\Sigma^{1/2}_{t}$, $s$)
21:     Reset random number generator with seed $s$triangleright For sampling $u_{i}$
22:     for $\theta_{i}\in\theta$ do
23:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
24:         $\theta_{i}\leftarrow\theta_{i}+\mu\Sigma^{1/2}_{t}u_{i}$triangleright Modify parameters in place
25:     end for
26:     return $\theta$
27:end function
Algorithm 1 HiZOO
We will present how to estimate Hessian inverse matrix $\Sigma$ in detail in Section 3.3. Given $\Sigma$, then we can construct the following descent direction:
$\begin{split}g_{\mu}(\theta_{t})=\sum\limits_{i=1}^{n}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma^{1/2}_{t}u_{i};\mathcal{B})-\mathcal{L}(\theta_{t}-\mu\Sigma^{1/2}_{t}u_{i};\mathcal{B})}{2\mu\cdot n}\cdot\Sigma_{t}^{1/2}u_{i}.\end{split}$ (1)
With the above descent direction, we can update $\theta_{t}$ as follows:
$\theta_{t+1}=\theta_{t}-\eta_{t}g_{\mu}(\theta_{t}).$ (2)
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1};\mathcal{B})\right]$ $=\mathcal{L}(\theta_{t};\mathcal{B})-\eta_{t}\mathbb{E}\left[\langle\nabla\mathcal{L}(\theta_{t};\mathcal{B}),g_{\mu}(\theta_{t})\rangle\right]+\mathcal{O}(\eta_{t}^{2})$
$=\mathcal{L}(\theta_{t};\mathcal{B})-\eta_{t}\frac{1}{b}\mathbb{E}\left[\sum_{i=1}^{b}\langle\nabla\mathcal{L}(\theta_{t};\mathcal{B}),\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B})\rangle\right]+\mathcal{O}(\eta_{t}^{2})+\mathcal{O}(\mu)$
$=\mathcal{L}(\theta_{t};\mathcal{B})-\eta_{t}\|\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|^{2}+\mathcal{O}(\eta_{t}^{2})+\mathcal{\mu},$
@@ 3.3 Diagonal Hessian Estimator
Storing the exact full spectral Hessian ($d\times d$) requires $\mathcal{O}(d^{2})$ memory [49, 48, 9], which is sufficient but never necessary. In HiZOO, we just estimate and retain only the diagonal Hessian which requires $\mathcal{O}(d)$ memory. It serves as a pre-conditioner to scale the direction and magnitude of the model parameter updates according to their respective curvatures.
Drawing from the lemma presented in MiNES [51]:
@@
where $H$ is the Hessian $\nabla^{2}\mathcal{L}(\theta;\mathcal{B})$ and $\Sigma$ is a positive definite matrix with $\Sigma^{-1}\approx\nabla^{2}\mathcal{L}(\theta;\mathcal{B})$.
Thus, we can approximate the diagonal Hessian by the zeroth order oracles. Firstly, we will access to the $\mathcal{L}(\theta+\mu\Sigma^{1/2}u;\mathcal{B})$, $\mathcal{L}(\theta-\mu\Sigma^{1/2}u;\mathcal{B})$ and $\mathcal{L}(\theta;\mathcal{B})$. Through the Taylor’s expansion, we yield the following results:
$\mathcal{L}(\theta+\mu\Sigma^{1/2}u;\mathcal{B})=\mathcal{L}(\theta;\mathcal{B})+\mu\langle\mathcal{L}(\theta;\mathcal{B}),\Sigma^{1/2}u\rangle+\frac{\mu^{2}}{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta;\mathcal{B})\Sigma^{1/2}u+\alpha(\theta,\mu\Sigma^{1/2}u).$ (4)
@@
$\mathcal{L}(\theta-\mu\Sigma^{1/2}u;\mathcal{B})=\mathcal{L}(\theta;\mathcal{B})-\mu\langle\mathcal{L}(\theta;\mathcal{B}),\Sigma^{1/2}u\rangle+\frac{\mu^{2}}{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta;\mathcal{B})\Sigma^{1/2}u+\alpha(\theta,-\mu\Sigma^{1/2}u).$
Then we can calculate the difference $\Delta\mathcal{L}$ by:
$\Delta\mathcal{L}$ $=\mathcal{L}(\theta+\mu\Sigma^{1/2}u;\mathcal{B})+\mathcal{L}(\theta-\mu\Sigma^{1/2}u;\mathcal{B})-2\mathcal{L}(\theta;\mathcal{B})$
$=\mu^{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta;\mathcal{B})\Sigma^{1/2}u+\alpha(\theta,\mu\Sigma^{1/2}u)+\alpha(\theta,-\mu\Sigma^{1/2}u).$
@@
$\frac{\Delta\mathcal{L}}{\mu^{2}}=u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta;\mathcal{B})\Sigma^{1/2}u+\mathcal{O}(\mu).$
@@
$\frac{1}{2}\mathbb{E}\bigg{[}\frac{\Delta\mathcal{L}}{\mu^{2}}\cdot\left(\Sigma^{-1/2}uu^{\top}\Sigma^{-1/2}-\Sigma^{-1}\right)\bigg{]}=\nabla^{2}\mathcal{L}(\theta;\mathcal{B})+\mathcal{O}(\mu).$
@@
Above equation shows that we can approximate the diagonal entries of $\nabla^{2}\mathcal{L}(\theta;\mathcal{B})$ by $\mathrm{diag}(\Sigma_{t}^{\prime}(\theta))$, requiring just one more forward pass per step compared with MeZO.
Due to the presence of noise in the calculation of the Hessian, we utilize exponential moving average (EMA) to denoise the diagonal Hessian estimation.
$\Sigma_{t+1}^{-1}=(1-\alpha_{t})\Sigma_{t}^{-1}+\alpha_{t}\left|\mathrm{diag}(\Sigma^{\prime}_{t})\right|.$ (6)